# Accelerating Clinical Evidence Synthesis with Large Language Models

### v1

@@ Abstract
Automatic medical discovery by AI is a dream of many. One step toward that goal is to create an AI model to understand clinical studies and synthesize clinical evidence from the literature. Clinical evidence synthesis currently relies on systematic reviews of clinical trials and retrospective analyses from medical literature. However, the rapid expansion of publications presents challenges in efficiently identifying, summarizing, and updating evidence. We introduce TrialMind, a generative AI-based pipeline for conducting medical systematic reviews, encompassing study search, screening, and data extraction phases. We utilize large language models (LLMs) to drive each pipeline component while incorporating human expert oversight to minimize errors. To facilitate evaluation, we also create a benchmark dataset TrialReviewBench, a custom dataset with 870 annotated clinical studies from 25 meta-analysis papers across various medical treatments. Our results demonstrate that TrialMind significantly improves the literature review process, achieving high recall rates (0.897-1.000) in study searching from over 20 million PubMed studies and outperforming traditional language model embeddings-based methods in screening (Recall@20 of 0.227-0.246 vs. 0.000-0.102). Furthermore, our approach surpasses direct GPT-4 performance in result extraction, with accuracy ranging from 0.65 to 0.84. We also support clinical evidence synthesis in forest plots, as validated by eight human annotators who preferred TrialMind over the GPT-4 baseline with a winning rate of 62.5%-100% across the involved reviews. Our findings suggest that an LLM-based clinical evidence synthesis approach, such as TrialMind, can enable reliable and high-quality clinical evidence synthesis to improve clinical research efficiency.
@@ Introduction
Clinical evidence is crucial for supporting clinical practices and advancing new drug development. It is primarily gathered through retrospective analysis of real-world data or through prospective clinical trials that assess new interventions on humans. Researchers often conduct systematic reviews to consolidate evidence from various studies in the literature [1]. However, conducting a systematic literature review is expensive and time-consuming, requiring five experts to analyze 195 publications over 67 weeks on average [2]. Moreover, the fast growth of clinical study databases means that the information in these published clinical reviews becomes outdated rapidly [3]. This situation underscores the urgent need to streamline the systematic review processes to produce systematic and timely clinical evidence from the extensive medical literature [4].
Large language models (LLMs) excel at information processing and generating, showing great promise in streamlining the clinical evidence synthesis process. Adapting LLMs to new tasks can be done by providing them the task definition and examples as the text inputs (namely ‘prompts’) without the need of retraining the model [5]. Researchers have tried to adopt LLMs for individual tasks in literature review [6, 7, 8, 9, 10]. For example, by enhancing inputs with multiple papers, LLMs can summarize findings to answer medical questions [9]. This strategy helps reduce hallucinations but still faces challenges when the input studies do not adequately support the posed questions, which requires more effort in literature search and screening steps. Furthermore, LLMs often demonstrate limitations in reasoning with numerical data found in clinical studies. The qualitative clinical evidence generated using raw paper content via prompting can be overly generic, lack critical information, or misinterpret results [11]. Therefore, we propose developing an LLM-driven pipeline to assist the entire workflow, including formulating research questions, conducting literature mining, extracting information, and synthesizing clinical evidence. This includes a comprehensive evaluation of LLMs in the entire process, which was not well-explored [12].
@@
Figure 1: The TrialMind framework has four components: a) Utilizing input PICO elements, TrialMind generates key terms to construct Boolean queries for retrieving studies from literature databases. b) TrialMind formulates eligibility criteria, which users can edit to provide context for LLMs during eligibility predictions. Users can then select studies based on these predictions and rank their relevance by aggregating them. c) TrialMind processes the descriptions of target data fields to extract and output the required information as structured data. d) TrialMind extracts findings from the studies and collaborates with users to synthesize the clinical evidence.
This study aims to fulfill the potential of AI in helping medical practitioners with the entire clinical evidence synthesis process using AI. We demonstrate how our method TrialMind is optimized for the clinical evidence synthesis task via: (1) generating boolean queries to search from the literature; (2) building inclusion and exclusion criteria to screen through the found studies; (3) extracting key information, including study protocols, methods, participant baselines, etc., from unstructured documents, per users’ request; and (4) synthesizing high-quality clinical evidence. Unlike previous approaches, TrialMind integrates LLMs into an AI pipeline by breaking down the task into multiple steps that align with expert practices. This approach maintains flexibility by involving humans in the loop to monitor, edit, and verify all intermediate steps and the final synthesized outputs.
@@
This paper curated a benchmark dataset TrialReviewBench for a comprehensive evaluation. The dataset includes 870 involved clinical studies and more than 50,000 identified studies from 25 meta-analyses. It also consists of manual annotations of 1,334 study characteristics and 1,049 study results. We show that the TrialMind is able to 1) retrieve a complete list of target studies from the literature, 2) follow the specified eligibility criteria to rank the most relevant studies at the top, and 3) achieve high accuracy in extracting information and clinical outcomes from unstructured documents based on user requests. Additionally, the extracted clinical outcomes can be standardized as input for meta-analysis (e.g., forest plots). We conducted a human evaluation of the synthesized evidence to demonstrate the potential value of TrialMind in practice.
@@ Results
A systematic understanding of cancer treatments is crucial for oncology drug discovery and development. Using a large list of cancer treatments from the National Cancer Institute (NCI) [13], we curated a dataset TrialReviewBench of systematic reviews. To ensure data quality, we crafted comprehensive queries with automatic filtering and manual screening. For each review, we obtained the list of studies with their PubMed IDs, retrieved their full content, and extracted study characteristics and clinical outcomes. We followed PubMed’s usage policy and guidelines during retrieval. Further manual checks were performed to correct inaccuracies, eliminate invalid and duplicate papers, and refine the text for clarity (Methods). The final TrialReviewBench dataset consists of 870 studies involved in 25 reviews (Fig. 2a), including (1) Brachytherapy [14, 15, 16]: 3 reviews with 189 involved studies and 5,811 identified studies; (2) Chemotherapy [17, 18, 19]: 3 reviews with 126 involved studies and 5,874 identified studies; (3) Hyperthermia [20, 21, 22]: 3 reviews with involved 28 studies and 5,972 identified studies; (4) Hormone Therapy [23, 24, 25]: 3 reviews with 30 involved studies and 5,970 identified studies; (5) Cancer Vaccine [26]: one review with 6 involved studies and 1,994 related studies; (6) Immune Checkpoint Inhibitors [27, 28, 29]: 3 reviews with 187 involved studies and 5,813 identified studies; (7) Immune System Modulators [30]: one review with 9 involved studies and 1,991 identified studies; (8) Monoclonal Antibodies [31, 32, 33]: 3 reviews with 228 involved studies and 5,772 identified studies; (9) T-cell Transfer Therapy [34, 35]: 2 reviews with 36 involved studies and 3,964 identified studies; (10) Stem Cell Transplant [36, 37, 38]: 3 reviews with 31 involved studies and 5,969 identified studies. The detailed characteristics of these studies are in Extended Table LABEL:tab:data_stats.
@@
Large language models (LLMs) excel in adapting to new tasks when provided with task-specific prompts while often struggling with complex tasks that require multi-step reasoning. Additionally, interacting and collaborating with LLMs can be problematic due to their opaque nature and the complexity of debugging [39]. In this study, we developed TrialMind that decomposes the clinical evidence synthesis process into four main tasks (Fig. 1 and Methods). Initially, using the provided research question enriched with population, intervention, comparison, and outcome (PICO) elements, TrialMind conducts a comprehensive search from the literature. It also works with users to build the criteria for study screening and rank the studies. Next, TrialMind browses the study details to extract the study characteristics and pertinent findings. To ensure the accuracy and integrity of the data, each output is linked to the sources for manual inspection. In the final step, TrialMind standardizes the clinical outcomes for meta-analysis.
Figure 2: Literature search experiment results. a, Total number of involved studies and participants of the review papers across different topics. b, The TrialMind’s interface for users to retrieve studies. c, The Recall of the search results for reviews across four topics. The bar heights indicate the Recall, and the star indicates the number of studies found. d, Scatter plots of the Recall against the number of ground-truth studies. Each scatter indicates the results of one review. Regression estimates are displayed with the 95% CIs in blue or purple.
@@
Finding relevant studies from medical literature like PubMed, which contains over 37 million entries, can be challenging. Typically, this requires the research expertise to craft complex queries that comprehensively cover pertinent studies. The challenge lies in balancing the specificity of queries: too stringent, and the search may miss relevant studies; too broad, and it becomes impractical to manually screen the overwhelming number of results. Previous approaches propose to prompt LLMs to generate the searching query in one pass [6], which can induce incomplete searching results due to the limited knowledge of LLMs. In contrast, TrialMind is designed to produce comprehensive queries through a pipeline that includes query generation, augmentation, and refinement. It also provides users with the ability to make adjustments (Fig. 2b).
@@
Overall, TrialMind achieved a Recall of 0.921 on average for all reviews in TrialReviewBench, meaning it can capture all target studies most of the time. By contrast, the GPT-4 baseline yielded Recall = 0.079, and the Human baseline yielded Recall = 0.230. We divided the search results across four topics determined by the treatments studied in each review (Fig. 2c). Our analysis showed that TrialMind can identify many more studies than the baselines. For instance, TrialMind achieved $\text{Recall}=0.914$ ($N$ studies = 28,863) for Immunotherapy-related reviews, while the GPT-4 baseline got $\text{Recall}=0.1$ ($N$ studies = 22), and the Human baseline got $\text{Recall}=0.154$ ($N$ studies = 153), respectively. In Radiation/Chemotherapy, TrialMind achieved $\text{Recall}=0.897$, the GPT-4 baseline got $\text{Recall}=0.017$, and the Human baseline got $\text{Recall}=0.304$. In Hormone Therapy, TrialMind achieved $\text{Recall}=0.924$, the GPT-4 baseline got $\text{Recall}=0.100$, and the Human baseline got $\text{Recall}=0.150$. In Hyperthermia, TrialMind achieved $\text{Recall}=1.000$, the GPT-4 baseline got $\text{Recall}=0.133$, and the Human baseline got $\text{Recall}=0.683$. These results demonstrate that regardless of the search task’s complexity, as indicated by the variability in the Human baseline, TrialMind consistently retrieves nearly all target studies from the PubMed database. This robust performance provides a solid foundation for accurately identifying target studies in the screening phase.
Furthermore, we made scatter plots of Recall versus the number of target studies for each review (Fig. 2d). The hypothesis was that an increase in target studies correlates with the difficulty of achieving complete coverage. Our findings reveal that TrialMind consistently maintained a high Recall, nearing 1 in many instances and never falling below 0.7, significantly outperforming the best baselines across all 25 reviews. A trend of declining Recall with an increasing number of target studies was confirmed through regression analysis. For example, with fewer than 20 target studies, TrialMind achieved perfect Recall for most reviews, while the GPT-4 baseline struggled, showing Recall close to 0, and the Human baseline results varied between 0 and 0.85. As the number of target studies increased, the Human and GPT-4 baselines’ Recall decreased to nearly zero. In contrast, TrialMind demonstrated remarkable resilience, showing minimal variation in performance despite the increasing number of target studies. For instance, in a review involving 141 studies, TrialMind achieved a Recall of 0.99, while the GPT-4 and Human baselines obtained a Recall of 0.02 and 0, respectively.
@@
[…]
Furthermore, TrialMind demonstrated significant improvements over the baselines across various therapeutic areas (Fig. 3b). For example, in “Cancer Vaccines” and “Hormone Therapy,” TrialMind substantially increased $\text{Recall}@50$, achieving 33.33-fold and 10.53-fold improvements, respectively, compared to the best-performing baseline. TrialMind generally attained a fold change greater than 2 (ranging from 1.57 to 33.33). Despite the challenge of selecting from a large pool of candidates ($n=2,000$) where candidates were very similar, TrialMind identified an average of 43% of target studies within the top 50.
We compared TrialMind to MedCPT and MPNet for $\text{Recall}@K$ ($K$ in 10 to 200) to gain insight into how $K$ influences the performances (Fig. 3e). We found TrialMind can capture most of the target studies when $K=200$, as it obtained $\text{Recall}@200=0.697$ in Immunotherapy, $\text{Recall}@200=0.618$ in Radiation/Chemotherapy, $\text{Recall}@200=0.744$ in Hormone Therapy, $\text{Recall}@200=0.830$ in Hyperthermia, respectively. The improvement over baselines is especially significant in Hormone Therapy and Hyperthermia, as the other baselines did not outperform the Random performances all the time. For instance, in Hormone Therapy, MedCPT’s $\text{Recall}@200=0.122$ and MPNet’s $\text{Recall}@200=0.083$ (Random’s $\text{Recall}@200=0.100$).
@@
Figure 5: Human evaluation of the synthesized clinical evidence. a, The study design compares the synthesized clinical evidence from the baseline and TrialMind via human evaluation. b, Winning rate of TrialMind against the GPT-4+Human baseline across studies. c, Violin plots of the ratings across studies. Each plot is tagged with the mean ratings (95% CI) from all the annotators. d,Violin plots of the ratings across annotators with different expertise levels. Each plot is tagged with the mean ratings (95% CI) from all the studies.
@@
TrialMind synthesizes clinical evidence from extracted results
We engaged human annotators to assess the quality of synthesized clinical evidence presented in forest plots, a format commonly used in systematic reviews to report meta-analysis results (Fig. 5a). We selected five systematic review studies as benchmarks and referenced the clinical evidence reported in the target studies. The baseline used GPT-4 with a simple prompting to extract the relevant text pieces that report the target outcome of interest (Methods). Manual calculations were necessary to standardize the data for meta-analysis. In contrast, TrialMind automated the extraction and standardization. Each annotator was asked to evaluate the evidence quality by comparing it against the evidence reported in the target review and deciding which method, TrialMind or the baseline, produced superior results. Additionally, they rated the quality of the synthesized clinical evidence on a scale of 1 to 5. The assignment of our method and the baseline was randomized to ensure objectivity.
The evaluation highlighted TrialMind’s superior performance compared to the direct use of GPT-4 for clinical evidence synthesis (Fig. 5b). We calculated the winning rate of TrialMind versus the baseline across the five studies. The results indicate a consistent preference by annotators for the evidence synthesized by TrialMind over that of the baseline. Specifically, TrialMind achieved winning rates of 87.5%, 100%, 62.5%, 62.5%, and 81.2%, respectively. The baseline’s primary shortcoming stemmed from the initial extraction step, where GPT-4 often failed to identify the relevant sources without well-crafted prompting. Therefore, the subsequent manual post-processing was unable to rectify these initial errors.
@@
Finally, we requested that annotators self-assess their expertise level in clinical studies, classifying themselves into three categories: ‘Basic’, ‘Familiar’, and ‘Advanced’. The typical profile ranges from computer scientists at the basic level to medical doctors at the advanced level. We then analyzed the ratings given to both methods across these varying expertise levels (Fig. 5d). We consistently observed higher ratings for TrialMind than the baseline across all groups. Annotators with basic knowledge tended to provide more conservative ratings, while those with more advanced expertise offered a wider range of evaluations. For instance, the ‘Basic’ group provided average ratings of 3.67 (95% CI 3.34-3.39) for TrialMind compared to 3.22 (95% CI 2.79-3.66) for the baseline. The ‘Advanced’ group rated TrialMind at an average of 3.40 (95% CI 3.16-3.64) and the baseline at 3.07 (95% CI 2.75-3.39).

### v2

@@ Abstract
Synthesizing clinical evidence largely relies on systematic reviews of clinical trials and retrospective analyses from medical literature. However, the rapid expansion of publications presents challenges in efficiently identifying, summarizing, and updating clinical evidence. Here, we introduce TrialMind, a generative artificial intelligence (AI) pipeline for facilitating human-AI collaboration in three crucial tasks for evidence synthesis: study search, screening, and data extraction. To assess its performance, we chose published systematic reviews to build the benchmark dataset, named TrialReviewBench, which contains 100 systematic reviews and the associated 2,220 clinical studies. Our results show that TrialMind excels across all three tasks. In study search, it generates diverse and comprehensive search queries to achieve high recall rates (Ours 0.711-0.834 v.s. Human baseline 0.138-0.232). For study screening, TrialMind surpasses traditional embedding-based methods by 30% to 160%. In data extraction, it outperforms a GPT-4 baseline by 29.6% to 61.5%. We further conducted user studies to confirm its practical utility. Compared to manual efforts, human-AI collaboration using TrialMind yielded a 71.4% recall lift and 44.2% time savings in study screening and a 23.5% accuracy lift and 63.4% time savings in data extraction. Additionally, when comparing synthesized clinical evidence presented in forest plots, medical experts favored TrialMind’s outputs over GPT-4’s outputs in 62.5% to 100% of cases. These findings show the promise of LLM-based approaches like TrialMind to accelerate clinical evidence synthesis via streamlining study search, screening, and data extraction from medical literature, with exceptional performance improvement when working with human experts.
@@ Introduction
Clinical evidence is crucial for supporting clinical practices and advancing new drug development and needs to be updated regularly [1]. It is primarily gathered through retrospective analysis of real-world data or through prospective clinical trials that assess new interventions on humans. Researchers usually conduct systematic reviews to consolidate evidence from various clinical studies in the literature [2, 3]. However, this process is expensive and time-consuming, requiring an average of five experts and 67.3 weeks based on an analysis of 195 systematic reviews [4]. Moreover, the fast growth of clinical study databases means that the information in these published clinical reviews becomes outdated rapidly [5]. For instance, PubMed has indexed over 35M citations and gets over 1M new citations annually [6]. This situation underscores the urgent need to streamline the systematic review processes to document systematic and timely clinical evidence from the extensive medical literature [7, 1].
Large language models (LLMs) excel at information processing and generating. They can be adapted to target tasks by providing the task definition and examples as the inputs (namely “prompts”) [8]. Researchers have tried to adopt LLMs for many individual tasks in the evidence synthesis process, including generating searching queries [9, 10], extracting studies’ population, intervention, comparison, outcome (PICO) elements [11, 12], screening citations [13], and summarizing findings from multiple studies [14, 15, 16, 17]. However, few have investigated LLMs’ effectiveness across the entire evidence synthesis process [18]. This is crucial because it ensures a seamless integration of AI in every step, potentially improving overall efficiency and accuracy. Understanding the strengths and limitations of LLMs in a holistic manner enables more effective automation and human-AI collaboration. To fill this gap, we created a testing dataset TrialReviewBench that covers major tasks in evidence synthesis, including study search, screening, and data extraction tasks. We chose published systematic reviews to create the dataset. As a result, the dataset includes 100 systematic reviews with 2,220 associated clinical studies. It also consists of manual annotations of 1,334 study characteristics and 1,049 study results. Based on TrialReviewBench, we are able to assess cutting-edge LLMs, e.g., GPT-4 [19], in clinical evidence synthesis tasks.
@@
Figure 1: The overview of TrialMind pipeline. a, it has four main steps: literature search, literature screening, data extraction, and evidence synthesis. b, (1) Utilizing input PICO elements, TrialMind generates key terms to construct Boolean queries for retrieving studies from literature databases. (2) TrialMind formulates eligibility criteria, which users can edit to provide context for LLMs during eligibility predictions. Users can then select studies based on these predictions and rank their relevance by aggregating them. (3) TrialMind processes the descriptions of target data fields to extract and output the required information as structured data. (4) TrialMind extracts findings from the studies and collaborates with users to synthesize the clinical evidence.
Furthermore, this study aims to fill the gap in adapting LLMs to evidence synthesis tasks, overcoming LLM’s limitations in (1) hallucinations, (2) weakness in reasoning with numerical data, (3) overly generic outputs, and (4) lack of transparency and reliability [20]. Specifically, we developed an AI-driven pipeline named TrialMind, which is optimized for (1) generating boolean queries to search citations from the literature; (2) building eligibility criteria and screening through the found citations; and (3) extracting data, including study protocols, methods, participant baselines, study results, etc., from publications and reports. More importantly, TrialMind breaks down into subtasks that adhere to the established practice of systematic reviews [21], which facilitates experts in the loop to monitor, edit, and verify intermediate outputs. It also has the flexibility to allow experts to begin at any intermediate step as needed.
@@
In this study, we show that the TrialMind is able to 1) retrieve a complete list of target studies from the literature, 2) follow the specified eligibility criteria to rank the most relevant studies at the top, and 3) achieve high accuracy in extracting information and clinical outcomes from unstructured documents based on user requests. Beyond providing descriptive evidence, TrialMind can extract numerical clinical outcomes to be standardized as input for meta-analysis (e.g., forest plots). A human evaluation was conducted to assess the synthesized evidence. Finally, to validate the practical benefits, we developed an accessible web application based on TrialMind and conducted a user study comparing two approaches: AI-assisted experts versus standalone experts. We measured the time savings and evaluated the output quality of each approach. The results show that TrialMind significantly reduced the time required for study search, citation screening, and data extraction, while maintaining or improving the quality of the output compared to experts working alone.
@@ Results
A systematic understanding of cancer treatments is crucial for oncology drug discovery and development. We retrieved a list of cancer treatments from the National Cancer Institute’s introductory page as the keywords to search medical systematic reviews [22]. To ensure data quality, we crafted comprehensive queries with automatic filtering and manual screening. For each review, we obtained the list of studies with their PubMed IDs, retrieved their full content, and extracted study characteristics and clinical outcomes. We followed PubMed’s usage policy and guidelines during retrieval. Further manual checks were performed to correct inaccuracies, eliminate invalid and duplicate papers, and refine the text for clarity (Methods). The final TrialReviewBench dataset consists of 2,220 studies involved in 100 reviews (Fig. 2a), covering four major topics: Immunotherapy, Radiation/Chemotherapy, Hormone Therapy, and Hyperthermia. We manually created three major evaluation tasks based on these reviews: study search, study screening, and data extraction.
The study search task begins with the PICO (Population, Intervention, Comparison, Outcome) elements extracted from the abstract of a systematic review, which serve as the formal definition of the research question. The model being tested is tasked with generating relevant keywords for the treatment and condition terms, as depicted in Fig. 2e. These keywords are then used to form Boolean queries, which are submitted to search citations in the PubMed database. The performance of the model is evaluated by checking whether the retrieved studies include those that were actually involved in the target systematic review. The recall rate is computed by measuring the proportion of actually involved studies identified through the search.
For the study screening task, the input consists of the PICO elements defined in the target systematic review. A candidate set of 2,000 citations is created by combining the actual studies included in the review with additional citations retrieved during the search but not included in the review. The model being tested ranks these citations based on the likelihood that each citation should be included in the systematic review. To assess the model’s performance, we compute Recall@$k$: the recall value indicating how many of the actual included studies appear in the top $k$ ranked candidates.
The data extraction task focuses on retrieving specific information from the input study documents. In this case, we extract Table 1 from each systematic review, which typically details study characteristics such as study design, population demographics, and outcome measurements. These characteristics are matched to the individual studies and manually verified, yielding 1,334 study characteristic annotations. Additionally, we extract individual study results from the review’s reported analysis, often presented in forest plots, capturing metrics such as overall response and event rates, resulting in 1,049 study result annotations. The model being tested is given a list of target data points to extract, and its output is evaluated by assessing the accuracy of the extracted information based on the annotated datasets.
Figure 2: Literature search experiment results. a, The total number of involved studies and the number of review papers across different topics. b, The TrialMind’s interface for users to retrieve studies. c, the Recall of the search results for reviews across four topics. The bar heights indicate the Recall, and the star indicates the number of studies found. d, Scatter plots of the Recall against the number of ground-truth studies. Each scatter indicates the results of one review. Regression estimates are displayed with the 95% CIs in blue or purple. e, Example cases comparing the outputs of three methods.
@@
Large language models (LLMs) excel in adapting to new tasks when provided with task-specific prompts while often struggling with complex tasks that require multiple steps of planning and reasoning. Additionally, interacting and collaborating with LLMs can be problematic due to their opaque nature and the complexity of debugging [23]. In this study, we developed TrialMind that decomposes the clinical evidence synthesis process into four main tasks (Fig. 1 and Methods). Initially, using the provided research question enriched with population, intervention, comparison, and outcome (PICO) elements, TrialMind conducts a comprehensive search from the literature. It also works with users to build the eligibility criteria for target studies and then automate screening and ranking identified citations. Next, TrialMind browses the study details to extract the study characteristics and pertinent findings. To ensure the accuracy and integrity of the data, each output is linked to the sources for manual inspection. In the final step, TrialMind standardizes the clinical outcomes for meta-analysis.
@@
Finding relevant studies from medical literature like PubMed, which contains over 35 million entries, can be challenging. Typically, this requires the research expertise to craft complex queries that comprehensively cover pertinent studies. The challenge lies in balancing the specificity of queries: too stringent, and the search may miss relevant studies; too broad, and it becomes impractical to manually screen the overwhelming number of results. Previous approaches propose to prompt LLMs to generate the searching query directly [9], which can induce incomplete searching results due to the limited knowledge of LLMs. In contrast, TrialMind is designed to produce comprehensive queries through a pipeline that includes query generation, augmentation, and refinement. It also provides users with the ability to make further adjustments (Fig. 2b).
@@
Overall, TrialMind achieved a Recall of 0.782 on average for all reviews in TrialReviewBench, meaning it can capture most of the target studies. By contrast, the GPT-4 baseline yielded Recall = 0.073, and the Human baseline yielded Recall = 0.187. We divided the search results across four topics determined by the treatments studied in each review (Fig. 2c). Our analysis showed that TrialMind can identify many more studies than the baselines. For instance, TrialMind achieved $\text{Recall}=0.797$ with identified studies $N$ = 22,084 for Immunotherapy-related reviews, while the GPT-4 baseline got $\text{Recall}=0.094$ ($N$ studies = 27), and the Human baseline got $\text{Recall}=0.154$ ($N$ studies = 958), respectively. In Radiation/Chemotherapy, TrialMind achieved $\text{Recall}=0.780$, the GPT-4 baseline got $\text{Recall}=0.020$, and the Human baseline got $\text{Recall}=0.138$. In Hormone Therapy, TrialMind achieved $\text{Recall}=0.711$, the GPT-4 baseline got $\text{Recall}=0.067$, and the Human baseline got $\text{Recall}=0.232$. In Hyperthermia, TrialMind achieved $\text{Recall}=0.834$, the GPT-4 baseline got $\text{Recall}=0.106$, and the Human baseline got $\text{Recall}=0.202$. These results demonstrate that regardless of the search task’s complexity, as indicated by the variability in the Human baseline, TrialMind consistently retrieves nearly all target studies from the PubMed database. This robust performance provides a solid foundation for accurately identifying target studies in the screening phase.
Furthermore, we made scatter plots of Recall versus the number of target studies for each review (Fig. 2d). The hypothesis was that an increase in target studies correlates with the difficulty of achieving complete coverage. Our findings reveal that TrialMind consistently maintained a high Recall, significantly outperforming the best baselines across all 100 reviews. A trend of declining Recall with an increasing number of target studies was confirmed through regression analysis. It was found that the GPT-4 baseline struggled, showing Recall close to 0, and the Human baseline results varied, with most reviews below 0.5. As the number of target studies increased, the Human and GPT-4 baselines’ Recall decreased to nearly zero. In contrast, TrialMind demonstrated remarkable resilience, showing minimal variation in performance despite the increasing number of target studies. For instance, in a review involving 141 studies, TrialMind achieved a Recall of 0.99, while the GPT-4 and Human baselines obtained a Recall of 0.02 and 0, respectively.
@@
We found that TrialMind greatly improved ranking performances, with the fold changes over the best baselines ranging from 1.3 to 2.6 across four topics (Table 3c). For instance, for the Hormone Therapy topic, TrialMind obtained $\text{Recall}@20=0.431$ and $\text{Recall}@50=0.674$. In the Hyperthermia topic, TrialMind obtained $\text{Recall}@20=0.518$ and $\text{Recall}@50=0.710$. In the Immunotherapy topic, TrialMind obtained $\text{Recall}@20=0.567$ and $\text{Recall}@50=0.713$. In the Radiation/Chemotherapy topic, TrialMind obtained $\text{Recall}@20=0.416$ and $\text{Recall}@50=0.654$. In contrast, other baselines exhibit significant variability across different topics. The general domain baseline MPNet was the worst as it performed similarly to the Random baseline in $\text{Recall}@20$. MedCPT showed marginal improvement over MPNet in the last three topics, while both failed to capture enough target studies in all topics.
Furthermore, TrialMind demonstrated significant improvements over the baselines across various therapeutic areas (Fig. 3b). For example, in “Cancer Vaccines” and “Hormone Therapy,” TrialMind substantially increased $\text{Recall}@50$, achieving 33.33-fold and 10.53-fold improvements, respectively, compared to the best-performing baseline. TrialMind generally attained a fold change greater than 2 (ranging from 1.57 to 33.33). Despite the challenge of selecting from a large pool of candidates ($n=2,000$) where candidates were very similar, TrialMind identified an average of 43% of target studies within the top 50. We compared TrialMind to MedCPT and MPNet for $\text{Recall}@K$ ($K$ in 10 to 200) to gain insight into how $K$ influences the performances (Fig. 3e). We found TrialMind can capture most of the target studies (over 80%) when $K=100$.
@@
Figure 5: Results of human evaluation and user study. a, TrialMind’s result extraction process. b, Winning rate of TrialMind against the GPT-4+Human baseline across studies. c, Violin plots of the ratings across studies. Each plot is tagged with the mean ratings (95% CI) from all the annotators. d,Violin plots of the ratings across annotators with different expertise levels. Each plot is tagged with the mean ratings (95% CI) from all the studies. e, Overall performance and time cost of study screening and data extraction tasks, respectively. f, Screening time cost and performance across reviews and participants. g, Data extraction accuracy across participants and different types of data.
@@
TrialMind facilitates clinical evidence synthesis via human-AI collaboration
We selected five systematic review studies as benchmarks and referenced the clinical evidence reported in the target studies. The baseline used GPT-4 with a simple prompting to extract the relevant text pieces that report the target outcome of interest (Methods). Manual calculations were necessary to standardize the data for meta-analysis. In contrast, TrialMind automated the extraction and standardization (Fig. 5a by (1) extracting the raw result description from the input document and (2) standardizing the results by generating a Python program to assist the calculation. The standardized results from all involved studies are then fed into the R program by human experts to make the aggregated evidence in a forest plot.
We engaged with human annotators to assess the quality of synthesized clinical evidence presented in forest plots. Each annotator was asked to evaluate the evidence quality by comparing it against the evidence reported in the target review and deciding which method, TrialMind or the baseline, produced superior results (Extended Fig. 1). Additionally, they rated the quality of the synthesized clinical evidence on a scale of 1 to 5. The assignment of our method and the baseline was randomized to ensure objectivity. The results highlighted TrialMind’s superior performance compared to the direct use of GPT-4 for clinical evidence synthesis (Fig. 5b). We calculated the winning rate of TrialMind versus the baseline across the five studies. The results indicate a consistent preference by annotators for the evidence synthesized by TrialMind over that of the baseline. Specifically, TrialMind achieved winning rates of 87.5%, 100%, 62.5%, 62.5%, and 81.2%, respectively. The baseline’s primary shortcoming stemmed from the initial extraction step, where GPT-4 often failed to identify the relevant sources without well-crafted prompting. Therefore, the subsequent manual post-processing was unable to rectify these initial errors.
@@
We requested that annotators self-assess their expertise level in clinical studies, classifying themselves into three categories: ‘Basic’, ‘Familiar’, and ‘Advanced’. The typical profile ranges from computer scientists at the basic level to medical doctors at the advanced level. We then analyzed the ratings given to both methods across these varying expertise levels (Fig. 5d). We consistently observed higher ratings for TrialMind than the baseline across all groups. Annotators with basic knowledge tended to provide more conservative ratings, while those with more advanced expertise offered a wider range of evaluations. For instance, the ‘Basic’ group provided average ratings of 3.67 (95% CI 3.34-3.39) for TrialMind compared to 3.22 (95% CI 2.79-3.66) for the baseline. The ‘Advanced’ group rated TrialMind at an average of 3.40 (95% CI 3.16-3.64) and the baseline at 3.07 (95% CI 2.75-3.39).
We conducted user studies to compare the quality and time efficiency between purely manual efforts and human-AI collaboration using TrialMind. Two participants were involved in both study screening and data extraction tasks. For the screening task, each participant was assigned 4 systematic review papers, with 100 candidate citations identified for each review. The participants were asked to select the 10 most likely relevant citations from the candidate pool. Each participant was provided with 2 candidate sets pre-ranked by TrialMind and 2 unranked sets. The participants also recorded the time taken to complete the screening process for each set. For the data extraction task, each participant was given 10 clinical studies. They manually extracted the target information for 5 of these studies. For the other 5, TrialMind was first used to perform an initial extraction, and the participants were required to verify and correct the extracted results. The time taken for the extraction process was reported for each study.
In Fig. 5e, we present the average performance and time cost for the AI+Human and Human-only approaches across both the study screening and data extraction tasks. The results demonstrate that the AI+Human approach consistently outperforms the Human-only approach. For the screening tasks, AI+Human achieved a 71.4% relative improvement in Recall, while reducing time by 44.2% compared to the Human-only arm. This underscores the significant advantage of TrialMind in accelerating the study screening process while also improving its quality. Similarly, for the data extraction tasks, the AI+Human approach improved extraction accuracy by 23.5% on average, with a 63.4% reduction in time required.
Detailed results of screening time and performance are shown in Fig. 5f, where two reviews showed the AI+Human approach achieving the same Recall as the Human-only arm with notable time savings, and in two other reviews, AI+Human achieved higher Recall with less time. From Fig. 5g, we see that the AI+Human approach delivered better or comparable accuracy across all three types of data, with the smallest gap in “Study design”. This is likely because study design information is often readily available in the study abstract, making it relatively easier for humans to extract. In contrast, the other two data types are embedded deeper within the main content, which can sometimes make it challenging for human readers to locate the correct information.